Use the determinant to determine whether the quadratic equation has two unequal real solutions, a double real solution or no real solution without solving the equation: x^2-2x+2=0
Answer by stanbon(75887)

x^2-2x+2=0
b^2-4ac=4-4(1)(2)<0
Therefore no Real Number solutions.
Cheers,
Stan H.